OpenStudy (anonymous):

FInd exact value of: csc (37) sec(53) - tan (53) cot (37)

OpenStudy (jdoe0001):

http://ichthyosapiens.com/School/Math/Trig/ch05CofunctionIdentities.jpg can you rewrite the csc(37) you think?

OpenStudy (jdoe0001):

keep in mind that 90-53 = 37

OpenStudy (anonymous):

Ok, I'll try, hold on...

OpenStudy (anonymous):

1/sin (37) * 1/cos (37) - sin (37)/cos (37) * cos (37)/sin (37) Is this right?

OpenStudy (jdoe0001):

hmm did you see the "cofunction identities" ?

OpenStudy (anonymous):

Oh, ok I'll look...

OpenStudy (anonymous):

1/sin (37) * 1/sin (37) - cos (37)/sin (37) * cos (37)/sin (37)

OpenStudy (jdoe0001):

hmm gimme one sec

OpenStudy (anonymous):

Ok.

OpenStudy (jdoe0001):

so using the cofunction identities -> http://ichthyosapiens.com/School/Math/Trig/ch05CofunctionIdentities.jpg then we get \( \large \begin{array}{llll} csc (37^o) &sec(53^o) - tan (53^o)& cot (37^o)\\ \quad \\ csc(90^o-{\color{red}{ 53}}^o)&sec(53^o) -tan (53^o)&cot(90^o-{\color{red}{ 53}}^o)\\ \quad \\ sec({\color{red}{ 53}}^o)&sec(53^o) -tan (53^o)&tan({\color{red}{ 53}}^o) \end{array}\)

OpenStudy (jdoe0001):

so... what would that give you anyway?

OpenStudy (anonymous):

I thought I was supposed to put everything in terms of sin and cos, then simply...?

OpenStudy (jdoe0001):

well... yes, I gather you could do it that way too

OpenStudy (anonymous):

So does that make my second answer (without the simplication) correct?

OpenStudy (anonymous):

Or did I do that wrong?

OpenStudy (jdoe0001):

nope, looks good the way you have it

OpenStudy (jdoe0001):

using the cofunctions identities you'd end up with -> 1/sin (37) * 1/sin (37) - cos (37)/sin (37) * cos (37)/sin (37)

OpenStudy (anonymous):

1/sin (37) * 1/sin (37) - cos (37)/sin (37) * cos (37)/sin (37) Simplified: 1/sin (37) - 2 cos (37)/sin (37) ? Not sure how to simplify this further...?

OpenStudy (anonymous):

Wouldn't that give me a negative? -cos (37)/sin (37)

OpenStudy (jdoe0001):

\(\bf \left[\cfrac{1}{sin(37^o)}\cdot \cfrac{1}{sin(37^o)}\right]-\left[\cfrac{cos(37^o)}{sin(37^o)}\cdot \cfrac{cos(37^o)}{sin(37^o)}\right]\\ \quad \\ \implies \cfrac{1^2}{sin^2(37^o)}-\cfrac{cos^2(37^o)}{sin^2(37^o)}\)

OpenStudy (anonymous):

or -tan (37) ?

OpenStudy (jdoe0001):

well, not quite

OpenStudy (jdoe0001):

if we add those fractions, we get \(\bf \cfrac{1^2}{sin^2(37^o)}-\cfrac{cos^2(37^o)}{sin^2(37^o)}\implies \cfrac{{\color{blue}{ 1-cos^2(37^o)}}}{sin^2(37^o)}\\ \quad \\ recall\implies sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)={\color{blue}{ 1-cos^2(\theta)}}\)

OpenStudy (anonymous):

I don't see how you got 1- cos^2 (37)...?

OpenStudy (jdoe0001):

well notice the LCD from the two fractions

OpenStudy (jdoe0001):

both fractions have the same denominator, thus the LCD will be just that, then the numerator will simply be itself moved over

OpenStudy (jdoe0001):

\(\bf 1^2=1\)

OpenStudy (anonymous):

Thanks.

OpenStudy (jdoe0001):

yw